# Exact Model Reduction for Continuous-Time Open Quantum Dynamics

### v2

@@ II.1 Dynamical settings and problem statement
with the reduced dynamics of $\rho_{S}(t)$ being non-unitary (and generally non-Markovian) even in the case where $\mathcal{D}_{L_{u}}\equiv 0$.
Regardless, in many situations of fundamental and practical relevance, interest may be a priori restricted to a subset of initial input states, and a subset of output quantities that depend upon the final, time-evolved state $\rho(t)$ and may be directly associated to or required for computing experimentally accessible properties. For instance, one may want to be able to predict:
@@
II.2 Reductions to operator subspaces
@@
and associating to each $\rho\in\mathfrak{S}$ a corresponding $\rho_{\cal X}={\cal R}(\rho).$ Of course, in general such a reduced model will not reproduce correctly the output of the initial model, nor will it maintain the requisite CPTP properties. The following sections focus on the construction of $\mathscr{W}$ and ${\cal X}$ so that the reduced model solve the problems we posed above.
III Reachable-based model reduction
@@
III.1 Minimal reachable linear models and their limitations
@@
Note that the reachable subspace $\mathscr{R}$ is effectively a Krylov operator subspace [39] and, from a system-theory viewpoint, it can be seen as the subspace reachable from the input of a fictitious linear input-output system (see e.g., [71, Sec. III.A]). As in the well-studied classical case, $\mathscr{R}$ enjoys some relevant properties that are summarized in the following proposition, whose proof can be found in Appendix A:
@@
The proof can also be found in Appendix A.
@@
A natural approach for using $\mathscr{R}$ to build an algebraic model is to consider the smallest *-algebra containing $\mathscr{R},$ $\mathscr{A}=\textrm{alg}(\mathscr{R}).$ This choice is particularly convenient, as any orthogonal projection onto a *-subalgebra $\Pi_{\mathscr{A}}$ can be shown to be a CPTP map [74, Thorem II.6.10.2], which admits a further factorization in full-rank maps of the form $\Pi_{\mathscr{A}}={\cal J}\circ{\cal R}$ as in Eq. (7), with all maps involved being also CPTP [49, Theorem 1]. This property, as we shall see in the following, is sufficient to ensure that the reduced generator ˇ$\check{\mathcal{L}}=\mathcal{R}\mathcal{L}\mathcal{J}$ is also a Lindblad generator. However, this choice of algebra is in general non-optimal, as illustrated by the following example.
@@
III.2 Algebraic model reduction: Fundamentals
III.2.1 Finite-dimensional *-algebras and block-representations
@@
A key result regarding the structure of *-algebras, which has been extensively used in the theory of quantum error correction and virtual quantum subsystems [62, 76, 77, 67, 68, 78], is the Wedderburn decomposition [79, 80]. Given any associative *-algebra $\mathscr{A}\subseteq\mathfrak{B}(\mathcal{H})$, there exists a decomposition of the Hilbert space, say,
@@
III.2.2 Quantum conditional expectations and distorted algebras
@@
Note that while idempotent, a conditional expectation need not be self-adjoint and thus, in general, it is not an orthogonal projection. The adjoint operator of the conditional expectation with respect to the Hilbert-Schmidt inner product, denoted by $\mathbb{J}|_{\mathscr{A}}(\cdot)\equiv\mathbb{E}|_{\mathscr{A}}^{\dagger}(\cdot)$, is known as the state extension [83]. Some properties of the conditional expectation directly follow from its definition. In particular, since $\mathscr{A}$ is unital, the conditional expectation $\mathbb{E}|_{\mathscr{A}}(\cdot)$ is a positive, unital projector. Then its dual, $\mathbb{J}|_{\mathscr{A}}(\cdot)$ is a positive and TP projector. It is possible to show that, in fact, both maps are actually CP [74].
The state extension is the map we need in order to reduce the description of the system in Schrödinger’s picture. Consider $X\in\mathscr{A}\subsetneq\mathfrak{B}(\mathcal{H})$ and let $\rho\in\mathfrak{D}(\mathcal{H}).$ Then, in computing the expectation for $X$ according to $\rho$ we have:
@@
It is straightforward to show that $\mathscr{A}_{\tau}\equiv\textrm{image}(\mathbb{J}|_{\mathscr{A}})$ is a linear operator subspace that is *-closed and closed with respect to the $\circ_{\tau}$ product. Sets of this type are called distorted algebras [75, 84] and are often used to characterize the set of fixed points of CPTP maps [85, 78].
@@
Consider an operator subspace $\mathscr{V}\subseteq\mathfrak{B}(\mathcal{H})$ and a positive-definite operator $V\in\mathscr{V}$, $V=V^{\dagger}>0$. Define
@@
III.2.3 Reduced representation of states and observables
@@
The Wedderburn decomposition of $\mathscr{A}$ let us naturally reduce the size of the representation of a *-subalgebra by removing the repeated blocks. Explicitly, if $\mathscr{A}=U\left(\bigoplus_{k}\mathfrak{B}(\mathcal{H}_{F,k})\otimes\mathbb{1}_{G,k}\right)U^{\dagger}$, with $m\equiv\sum_{k}\dim(\mathcal{H}_{F,k})$, we can observe that it is isomorphic to $\check{\mathscr{A}}=\bigoplus_{k}\mathfrak{B}(\mathcal{H}_{F,k})\subseteq\mathbb{C}^{m\times m}$ (see Fig. 1 for a pictorial representation). If we aim to reduce a model onto a *-sub-algebra $\mathscr{A}$, it then suffices to consider its compressed representation $\check{\mathscr{A}}$. As shown in [49], this reduction in representation can be achieved by using CP unital or CPTP maps that factorize the conditional expectation or its dual, respectively:
@@
Figure 1: Graphical representation of the operator *-algebra $\mathscr{A}=(\mathbb{1}_{2}\otimes\mathbb{C}^{4\times 4})\bigoplus(\mathbb{1}_{3}\otimes\mathbb{C}^{2\times 2})\bigoplus\mathbb{C}^{3\times 3}\subseteq\mathbb{C}^{17\times 17}$ and its compressed representation $\check{\mathscr{A}}=\mathbb{C}^{4\times 4}\bigoplus\mathbb{C}^{2\times 2}\bigoplus\mathbb{C}^{3\times 3}\subseteq\mathbb{C}^{9\times 9}$. The action of the reduction and injection maps $\mathcal{R}$ and $\mathcal{J}$ is also shown.
@@
III.3 Reduction of semigroup dynamics on *-subalgebras
Before summarizing our MR procedure, we need to determine what happens to the restriction of a Lindblad generator $\mathcal{L}$ to a distorted algebra. While, for discrete-time dynamics, the reduction automatically preserves the CPTP character of the original map since $\mathcal{J},\mathcal{R}$ are also CPTP, in the continuous-time case it is not a priori obvious that the reduction yields a valid Lindblad generator. In the following theorem, which is one of our main results, we prove that the reduction is indeed a CPTP semigroup generator that leaves the algebra invariant. While some supporting mathematical results are provided in Appendix A, the main idea of the proof is simple and we include it here due to the importance of the result.
@@ III.3 Reduction of semigroup dynamics on *-subalgebras
From Proposition 4 in Appendix A, we then have that ˇ$\check{\mathcal{L}}(\check{\rho})=\mathcal{R}\Phi\mathcal{J}(\check{\rho})-\mathcal{J}^{\dagger}(K)\check{\rho}-\check{\rho}\mathcal{J}^{\dagger}(K^{\dagger})$, where $\mathcal{J}^{\dagger}(K)\in\check{\mathscr{A}}$. Moreover, from Throrem 3, we know that both $\mathcal{R}$ and $\mathcal{J}$ are CPTP and hence $\mathcal{R}\circ\Phi\circ\mathcal{J}$ is CP. To conclude, we can verify that ˇ$\check{\mathcal{L}}^{\dagger}(\mathbb{1})=0$ by observing that
@@
IV Observable-based model reduction
IV.1 Minimal observable linear models
In this section we tackle a MR problem that is dual to the one considered in Sec. III, namely, we assume no restriction on the initial conditions, $\mathfrak{S}=\mathfrak{D}(\mathcal{H})$. In this case, the reduction can only leverage the observables of interest: intuitively, the key idea here is to discard from the model degrees of freedom that produce no output.
@@ IV.1 Minimal observable linear models
IV.2 Extension to quantum models
@@
    Compute [^1] It is worth noting that, to the purpose of obtaining the relevant Wedderburn decomposition, computing the commutant $(\mathscr{N}^{\perp})^{\prime}$ is equivalent, since the Wedderburn decomposition of the commutant automatically gives the one of the algebra, see e.g., Eqs. (12) and (13). Interestingly, in implementing this algorithm for the examples of Sec. VI, we observed that the numerical calculations of the Wedderburn decomposition were more efficient and numerically stable through the calculation of the commutant instead of the algebra itself. the output algebra $\mathscr{O}={\rm alg}(\mathscr{N}^{\perp}).$ This can be done as a double commutant $\mathscr{O}=(\mathscr{N}^{\perp})^{\prime\prime}.$ Given $\mathscr{O}$, find the unitary change of base $U$ that brings $\mathscr{N}^{\perp}$ and $(\mathscr{N}^{\perp})^{\prime}$ to their canonical Wedderburn decomposition.
@@
Notice that in the above procedure for observable-based MR, differently from the reachable-based case, we close $\mathscr{N}^{\perp}$ to a standard algebra, rather than a distorted one. One may wonder whether, by using a distorted algebra, one could obtain a smaller reduced model also in the observable MR; however, this is not the case. Intuitively, one can understand this as follows: algebras are images of conditional expectations, which are maps acting on observables; as such, they form the natural spaces to describe observables. Distorted algebras are images of state extensions, which act on states; hence they are the natural structure to describe states. As shown in Eq. (25), $\mathscr{N}^{\perp}$ is a space of Heisenberg-evolved observables and it is thus natural to close this space to an algebra, not to a distorted one. Formally proving that there is no advantage in closing $\mathscr{N}^{\perp}$ to a distorted algebra is rather involved, and we do not further elaborate on that here (see [49, Theorem 6] for technical details).
IV.3 Joint reachable- and observable-based model reduction
@@
V Symmetry-based model reduction
@@
V.1 Symmetries for quantum dynamical semigroups
@@
V.2 Symmetries and invariant subspaces
@@
V.3 Symmetries and observable-based model reduction
@@
since ${\rm alg}\{\mathscr{N}^{\perp}\}$ is the smallest *-algebra that contains $\mathscr{N}^{\perp}$ by definition. This allows us to conclude that: (i) The algebra $\mathbb{C}\mathscr{G}^{\prime}$, via Theorem 6 and the procedure described in Sec. IV.2, may be used to obtain a valid quantum MR, with $\mathscr{O}\equiv\mathbb{C}\mathscr{G}^{\prime}$. (ii) The resulting observable-based MR is (at least) as good as a reduction that directly exploits the knowledge of symmetries, with the associated irrep block-decomposition of Eqs. (30)-(31) being computed through known algorithms [81, 96, 97, 43]. Note that knowing that a model exhibits some symmetries may help in computing $\mathscr{N}^{\perp}$, as one could perform the symmetry-based MR first, then compute $\mathscr{N}^{\perp}$ on the resulting reduced model more efficiently and verify whether further reduction is possible or not.
@@
V.4 Observable-dependent symmetries
@@
The proof follows identical steps to those in Proposition 3. Remarkably, as we formally prove in Appendix A, reducing the dynamics to the commutant of the ODS group of interest turns out to be equivalent to constructing the reduction to the output algebra $\mathscr{O}$:
@@ VI Illustrative applications
As usual, $\sigma_{q},$ with $q\in\{0,x,y,z\}$ and $\sigma_{0}\equiv\mathbb{1}_{2}$, denote Pauli spin-1/2 matrices, while we will use the shorthand notation $\sigma_{q}^{(k)}\equiv\otimes_{i=1}^{k-1}\sigma_{0}\otimes\sigma_{q}\otimes_{i=1}^{N-k}\sigma_{0}$ to denote their multi-spin extension, with $\sigma_{q}^{(k)}\in\mathfrak{B}(\mathcal{H})$, acting non-trivially only on the $k$-th spin when $q\equiv u\in\{x,y,z\}$. Similarly, we will define local ladder (raising and lowering) operators acting non-trivially on spin $k$ as $\sigma_{\pm}^{(k)}\equiv\frac{1}{2}(\sigma_{x}^{(k)}\pm i\sigma_{y}^{(k)})$.
VI.1 Dissipative central spin model

### v3

@@ 2.1 Dynamical settings and problem statement
with the reduced dynamics of $\rho_{S}(t)$ being non-unitary (and generally non-Markovian) even in the case where $\mathcal{D}_{L_{u}}\equiv 0$. Regardless, in many situations of fundamental and practical relevance, interest may be a priori restricted to a subset of initial input states, and a subset of output quantities that depend upon the final, time-evolved state $\rho(t)$ and may be directly associated to or required for computing experimentally accessible properties. For instance, one may want to be able to predict:
@@
2.2 Reductions to operator subspaces
@@
and associating to each $\rho\in\mathfrak{S}$ a corresponding $\rho_{\cal X}={\cal R}(\rho).$ Of course, in general such a reduced model will not reproduce correctly the output of the initial model, nor will it maintain the requisite CPTP properties. The following sections focus on constructing $\mathscr{W}$ and ${\cal X}$ so that the resulting reduced model does solve the problems we posed above.
3 Reachable-based model reduction
@@
3.1 Minimal reachable linear models and their limitations
@@
Note that the reachable subspace $\mathscr{R}$ is effectively a Krylov operator subspace [39] and, from a system-theory viewpoint, it can be seen as the subspace reachable from the input of a fictitious linear input-output system (see e.g., [75, Sec. III.A]). As in the well-studied classical case, $\mathscr{R}$ enjoys some relevant properties that are summarized in the following proposition, whose proof can be found in Appendix A.1:
@@
The proof can also be found in Appendix A.1.
@@
A natural approach for using $\mathscr{R}$ to build an algebraic model is to consider the smallest *-algebra containing $\mathscr{R},$ $\mathscr{A}=\textrm{alg}(\mathscr{R}).$ This choice is particularly convenient, as any orthogonal projection onto a *-subalgebra $\Pi_{\mathscr{A}}$ can be shown to be a CPTP map [78, Thorem II.6.10.2], which admits a further factorization in full-rank maps of the form $\Pi_{\mathscr{A}}={\cal J}\circ{\cal R}$ as in Eq. (7), with all maps involved being also CPTP [51, Theorem 1]. This property, as we shall see in the following, is sufficient to ensure that the reduced generator ˇ$\check{\mathcal{L}}=\mathcal{R}\mathcal{L}\mathcal{J}$ is also a Lindblad generator. However, this choice of $\mathscr{A}$ is in general non-optimal, as illustrated by the following example.
@@
3.2 Algebraic model reduction: Fundamentals
3.2.1 Finite-dimensional *-algebras and their representations
@@
A key result regarding the structure of *-algebras, which has been extensively used in the theory of virtual quantum subsystems [66, 80, 81] and information-preserving structures [71],[79],[72], is the Wedderburn decomposition [62, 63]. Given any associative *-algebra $\mathscr{A}\subseteq\mathfrak{B}(\mathcal{H})$, there exists a decomposition of the Hilbert space, say,
@@
3.2.2 Quantum conditional expectations and distorted algebras
@@
While idempotent, a conditional expectation need not be self-adjoint and thus, in general, it is not an orthogonal projection. The adjoint operator of the conditional expectation with respect to the Hilbert-Schmidt inner product, denoted by $\mathbb{J}|_{\mathscr{A}}(\cdot)\equiv\mathbb{E}|_{\mathscr{A}}^{\dagger}(\cdot)$, is known as the state extension [84]. Some properties of the conditional expectation directly follow from its definition. In particular, since $\mathscr{A}$ is unital, the conditional expectation $\mathbb{E}|_{\mathscr{A}}(\cdot)$ is a positive, unital projector. Then its dual, $\mathbb{J}|_{\mathscr{A}}(\cdot)$ is a positive and TP projector. It is possible to show that, in fact, both maps are actually CP [78].
The state extension is the map we need to reduce the description of the system in the Schrödinger’s picture. Let $X\in\mathscr{A}\subsetneq\mathfrak{B}(\mathcal{H})$ and $\rho\in\mathfrak{D}(\mathcal{H}).$ Then, in computing the expectation for $X$ according to $\rho$ we have:
@@
It is easy to show that $\mathscr{A}_{\tau}\equiv\textrm{image}(\mathbb{J}|_{\mathscr{A}})$ is a linear operator subspace that is both *-closed and closed with respect to the $\circ_{\tau}$ product. Sets of this type are called distorted algebras [79, 85] and are often used to characterize the set of fixed points of CPTP maps [86, 87].
@@
Consider an operator subspace $\mathscr{V}\subseteq\mathfrak{B}(\mathcal{H})$ and a positive-definite operator $V\in\mathscr{V}$, $V=V^{\dagger}>0$. Let us define
@@
3.2.3 Reduced representation of states and observables
@@
As we anticipated in Sec. 1.2, the Wedderburn decomposition of $\mathscr{A}$ let us naturally reduce the size of the representation of a *-subalgebra by removing the repeated blocks. Explicitly, if $\mathscr{A}=U\left(\bigoplus_{k}\mathfrak{B}(\mathcal{H}_{F,k})\otimes\mathbb{1}_{G,k}\right)U^{\dagger},$ with $m\equiv\sum_{k}\dim(\mathcal{H}_{F,k})$, we can observe that it is isomorphic to $\check{\mathscr{A}}=\bigoplus_{k}\mathfrak{B}(\mathcal{H}_{F,k})\subseteq\mathbb{C}^{m\times m}$ (recall Fig. 1). If we aim to reduce a model onto a *-sub-algebra $\mathscr{A}$, it then suffices to consider its compressed representation $\check{\mathscr{A}}$. As shown in [51], such a reduction can be achieved by using CP unital or CPTP maps that factorize the conditional expectation or its dual, respectively:
@@
3.3 Reduction of continuous-time semigroup dynamics on *-subalgebras
Before summarizing our MR procedure, we need to determine what happens to the restriction of a Lindblad generator $\mathcal{L}$ to a distorted algebra. While, for discrete-time dynamics, the reduction automatically preserves the CPTP character of the original map since $\mathcal{J},\mathcal{R}$ are also CPTP, in the continuous-time case it is not a priori obvious that the reduction yields a valid Lindblad generator. In the following theorem, which is one of our main results, we prove that the reduction is indeed a CPTP semigroup generator that leaves the algebra invariant. While some supporting mathematical results are provided in Appendix A.1, the main idea of the proof is simple and we include it here due to the importance of the result.
@@ 3.3 Reduction of continuous-time semigroup dynamics on *-subalgebras
From Proposition 4 in Appendix A.1, we then have that ˇ$\check{\mathcal{L}}(\check{\rho})=\mathcal{R}\Phi\mathcal{J}(\check{\rho})-\mathcal{J}^{\dagger}(K)\check{\rho}-\check{\rho}\mathcal{J}^{\dagger}(K^{\dagger})$, where $\mathcal{J}^{\dagger}(K)\in\check{\mathscr{A}}$. Moreover, from Throrem 3, we know that both $\mathcal{R}$ and $\mathcal{J}$ are CPTP and hence $\mathcal{R}\circ\Phi\circ\mathcal{J}$ is CP. To conclude, we can verify that ˇ$\check{\mathcal{L}}^{\dagger}(\mathbb{1})=0$ by observing that
@@
4 Observable-based model reduction
4.1 Minimal observable linear models
In this section we tackle a MR problem that is dual to the one considered in Sec. 3, namely, we assume no restriction on the initial conditions, $\mathfrak{S}=\mathfrak{D}(\mathcal{H})$. In this case, the reduction can only leverage the observables of interest: intuitively, the key idea here is to discard from the model degrees of freedom that produce no output.
@@ 4.1 Minimal observable linear models
4.2 Extension to quantum models
@@
    Compute [^1] It is worth noting that, to the purpose of obtaining the relevant Wedderburn decomposition, computing the commutant $(\mathscr{N}^{\perp})^{\prime}$ is equivalent, since the Wedderburn decomposition of the commutant automatically gives the one of the algebra, see e.g., Eqs. (12) and (13). Interestingly, in implementing this algorithm for the examples of Sec. 6, we observed that the numerical calculations of the Wedderburn decomposition were more efficient and numerically stable through the calculation of the commutant instead of the algebra itself. the output algebra $\mathscr{O}={\rm alg}(\mathscr{N}^{\perp}).$ This can be done as a double commutant $\mathscr{O}=(\mathscr{N}^{\perp})^{\prime\prime}.$ Given $\mathscr{O}$, find the unitary change of base $U$ that brings $\mathscr{N}^{\perp}$ and $(\mathscr{N}^{\perp})^{\prime}$ to their canonical Wedderburn decomposition.
@@
Notice that in the above procedure for observable-based MR, differently from the reachable-based case, we close $\mathscr{N}^{\perp}$ to a standard algebra, rather than a distorted one. One may wonder whether, by using a distorted algebra, one could obtain a smaller reduced model also in the observable MR; however, this is not the case. Intuitively, one can understand this as follows: algebras are images of conditional expectations, which are maps acting on observables; as such, they form the natural spaces to describe observables. Distorted algebras are images of state extensions, which act on states; hence they are the natural structure to describe states. As shown in Eq. (25), $\mathscr{N}^{\perp}$ is a space of Heisenberg-evolved observables and it is thus natural to close this space to an algebra, not to a distorted one. Formally proving that there is no advantage in closing $\mathscr{N}^{\perp}$ to a distorted algebra is rather involved, and we do not further elaborate on that here (see [51, Theorem 6] for technical details).
4.3 Joint reachable- and observable-based model reduction
@@
5 Symmetry-based model reduction
@@
5.1 Symmetries for quantum dynamical semigroups
@@
5.2 Symmetries and invariant subspaces
@@
5.3 Symmetries and observable-based model reduction
@@
since ${\rm alg}\{\mathscr{N}^{\perp}\}$ is the smallest *-algebra that contains $\mathscr{N}^{\perp}$ by definition. This allows us to conclude that: (i) The algebra $\mathbb{C}\mathscr{G}^{\prime}$, via Theorem 6 and the procedure described in Sec. 4.2, may be used to obtain a valid quantum MR, with $\mathscr{O}\equiv\mathbb{C}\mathscr{G}^{\prime}$. (ii) The resulting observable-based MR is (at least) as good as a reduction that directly exploits the knowledge of symmetries, with the associated irrep block-decomposition of Eqs. (30)-(31) being computed through known algorithms [82, 98, 99, 43]. Note that knowing that a model exhibits some symmetries may help in computing $\mathscr{N}^{\perp}$, as one could perform the symmetry-based MR first, then compute $\mathscr{N}^{\perp}$ on the resulting reduced model more efficiently and verify whether further reduction is possible or not.
@@
5.4 Observable-dependent symmetries
@@
The proof follows identical steps to those in Proposition 3. Remarkably, as we formally prove in Appendix A.1, reducing the dynamics to the commutant of the ODS group of interest turns out to be equivalent to constructing the reduction to the output algebra $\mathscr{O}$:
@@ 6 Illustrative applications
As usual, $\sigma_{q},$ with $q\in\{0,x,y,z\}$ and $\sigma_{0}\equiv\mathbb{1}_{2}$, denote Pauli spin-1/2 matrices, while we will use the shorthand notation $\sigma_{q}^{(k)}\equiv\otimes_{i=1}^{k-1}\sigma_{0}\otimes\sigma_{q}\otimes_{i=k+1}^{N}\sigma_{0}$ to denote their multi-spin extension, with $\sigma_{q}^{(k)}\in\mathfrak{B}(\mathcal{H})$, acting non-trivially only on the $k$-th spin when $q\equiv u\in\{x,y,z\}$. Similarly, we will define local ladder (raising and lowering) operators acting non-trivially on spin $k$ as $\sigma_{\pm}^{(k)}\equiv\frac{1}{2}(\sigma_{x}^{(k)}\pm i\sigma_{y}^{(k)})$.
6.1 Dissipative central spin model